# Predicting solvation free energies with an implicit solvent machine learning potential

### v1

@@ Abstract
Machine learning (ML) potentials are a powerful tool in molecular modeling, enabling ab initio accuracy for comparably small computational costs. Nevertheless, all-atom simulations employing best-performing graph neural network architectures are still too expensive for applications requiring extensive sampling, such as free energy computations. Implicit solvent models could provide the necessary speed-up due to reduced degrees of freedom and faster dynamics. Here, we introduce a Solvation Free Energy Path Reweighting (ReSolv) framework to parametrize an implicit solvent ML potential for small organic molecules that accurately predicts the hydration free energy, an essential parameter in drug design and pollutant modeling. With a combination of top-down (experimental hydration free energy data) and bottom-up (ab initio data of molecules in a vacuum) learning, ReSolv bypasses the need for intractable ab initio data of molecules in explicit bulk solvent and does not have to resort to less accurate data-generating models. On the FreeSolv dataset, ReSolv achieves a mean absolute error close to average experimental uncertainty, significantly outperforming standard explicit solvent force fields. The presented framework paves the way toward deep molecular models that are more accurate yet computationally cheaper than classical atomistic models.
Introduction
@@
Solvation free energy, and notably hydration free energy, is generally recognized as a fundamental thermodynamic quantity of interest in computational chemistry. Defined as the work done when transferring a molecule from the gas phase to the solution (water in the case of hydration free energy), it enables the computation of several key physicochemical properties of molecules, such as solubility, partition coefficients, activity coefficients, and binding free energies in solutions [1, 2]. These properties are of great importance to the pharmaceutical, environmental, and materials sciences [3, 4, 5, 6, 7, 8, 9], prompting the organization of blind prediction SAMPL challenges [10, 11, 12] with hydration free energy as one of the main targets. In addition, Mobley et al. compiled and curated a FreeSolv database of experimentally measured hydration free energies for small neutral molecules in water [13, 14].
@@
A wide spectrum of methods is available to calculate solvation free energy, ranging from traditional approaches such as continuum solvation models [15, 16] to recent machine learning (ML) algorithms [17, 18, 19, 20, 21, 22, 23, 24, 25, 26] and their combinations [27, 28, 29]. The alchemical methods with Molecular Dynamics (MD) simulations [14, 30, 31] are typically assumed to be highly accurate but computationally expensive [32, 33]. However, both the fidelity and the efficiency highly depend on the explicitly treated degrees of freedom and the employed potential energy model.
@@
In implicit solvent models, the solvent molecules (e.g., water molecules) are not explicitly present in the system (as in explicit solvent models); instead, the interactions are modified to account for the solvent effects [34]. The number of degrees of freedom is thereby greatly reduced, resulting in large computational gains. Classical implicit models treat the solvent as a continuum medium with specific dielectric and interfacial properties. The typical approach decouples the electrostatic (polar) and nonpolar interactions. The former can be approximated by solving the Poisson–Boltzmann equation or further simplified with the popular generalized Born model, while the latter is most often estimated via the solvent-accessible surface area [34, 35]. While recent advances, such as using ML to predict the generalized Born radii [36], have increased the accuracy of these models, they are still in considerable disagreement with experimental data and explicit solvent models [37, 26, 38]. In particular, the solvation free energy root mean square error (RMSE) is approximately 3.6 kcal/mol for both the Poisson-Boltzmann and Generalized Born models. By specifically optimizing the nonpolar interactions using hydration free energy data, the test set RMSE can be decreased to 1.68 kcal/mol [39].
@@
Higher accuracy can be achieved if the two-body interaction approximation and ad-hoc modeling are replaced with a many-body, flexible potential energy surface, as is the case for ML potentials. During the past decade, ML potentials were deployed to derive atomistic [40] and coarse-grained [41] models based on training data provided by high-fidelity simulations [42], experiments [43], or both [44]. In the context of implicit water models, the capacity of deep ML potentials is large enough to compensate for the removal of solvent degrees of freedom. The predicted solute properties, such as the conformational landscape of proteins, match those obtained from the reference explicit water model [45, 46, 26, 47, 48, 49, 50, 51]. However, these studies used classical atomistic force fields as the data-generating model, inherently limiting the attainable accuracy of the resulting ML potential. As shown previously, and also in this work, classical atomistic models such as General Amber Force Field (GAFF) and CHARMM General Force Field (CGenFF) systematically overestimate hydration free energies [30]. Modifying the Lennard-Jones parameters yields an improved agreement with experiments but can, at the same time, negatively affect other properties [30].
@@
On the other hand, more accurate ab initio training data is prohibitively expensive for solutes in explicit bulk water. Consequently, frequently used Density Functional Theory (DFT) training databases such as QM9 [52], ANI [53], and QM7-X [54] contain small organic molecules in vacuum or at best contain samples with a few randomly placed water molecules around solutes [55, 56]. Moreover, implicit solvent models are coarse-grained models. When trained with the common force-matching approach [57], coarse-grained ML models require much more data than their atomistic counterparts [58]. The underlying reason is the surjective atomistic-to-coarse-grained mapping, resulting in noisy force labels. As an illustration, parametrization of coarse-grained models for alanine dipeptide in implicit water required a dataset with $10^{6}$ configurations [59, 42], a number easily obtainable with classical force fields but vastly out of reach for ab initio calculations.
@@
Results
Solvation Free Energy Path Reweighting (ReSolv)
@@ Results
The top-down training on experimental data is not as straightforward as bottom-up training in the first stage because $\Delta A$ is not an output of the ML model but rather evaluated from an MD simulation driven by an ML potential. We employ a variation of the Differential Trajectory Reweighting (DiffTRe) method [43] that avoids exploding gradients and reduces the computational and memory requirements compared to alternative gradient computation via backpropagation through the MD simulation. Before continuing, we first give a brief summary of the DiffTRe method.
Consider the task of matching a time-independent observable $O$, e.g., by postulating the following loss function $L=(\langle O\rangle_{\theta}-O_{\mathrm{exp}})^{2}$, where $\langle\rangle_{\theta}$ denotes the ensemble average with respect to the canonical distribution using ML potential $U_{\theta}=U(S;\theta)$. DiffTRe leverages thermodynamic perturbation theory [60], stating that $\langle O\rangle_{\theta}$ can be estimated from $N$ decorrelated states $\{S_{j}\}_{\hat{\theta}}$ generated by a reference potential $U_{\hat{\theta}}=U(S;\hat{\theta})$, that is
@@
where $\beta=1/(k_{B}T)$, $k_{B}$ is Boltzmann’s constant and $T$ temperature. Due to limited sampling, the estimation should only be used if the states generated by the reference potential $U_{\hat{\theta}}$ are statistically close to states that would have been sampled from the potential $U_{\theta}$. The distribution overlap is captured with the effective sample size [61] given by
@@
Here, we instead use the DiffTRe learning process itself as a free energy integration path. We initialize the parameters with the pre-trained $\theta_{\mathrm{vac}}$ parameters and iteratively perturb them during training towards the desired $\theta_{\mathrm{sol}}$ parameters (see Fig. 1). At each update step, we compute the free energy difference between the new and previous potential using a hybrid of the free energy perturbation [60] and the Bennett Acceptance Ratio (BAR) methods [62]. The free energy differences are accumulated along the training, which, at the end of the training, yield the total free energy difference between $U_{\mathrm{vac}}$ and $U_{\mathrm{sol}}$ or $\Delta A$.
Explicitly, let us consider the update step from $\theta_{i}$ to $\theta_{i+1}$. At this point, due to reweighting, the reference trajectory may have been generated at an earlier step $i-k$ with the ML potential $U(S;\theta_{i-k})$, where $k\leq i$. Since this is the reference trajectory, $\hat{\theta}=\theta_{i-k}$. As above, we denote the states generated by the reference potential with $\{S_{j}\}_{\hat{\theta}}$. The corresponding free energy difference is $\Delta A=\Delta A_{\theta_{0}\to\theta_{i-k}}$, which was already computed at step $i-k$. Firstly, we compute $N_{\mathrm{eff}}$ (Eq. 3). If the reweighting criterion is satisfied, the trajectory is reused. In the opposite case, we generate a new reference trajectory with the current ML potential $U(S;\theta_{i})$. Thus, $\hat{\theta}=\theta_{i}$. We also update $\Delta A=\Delta A+\Delta A^{BAR}_{\theta_{i-k}\to\theta_{i}}$, where $\Delta A^{BAR}_{\theta_{i-k}\to\theta_{i}}$ denotes the free energy difference between steps $i-k$ and $i$ estimated with the BAR method [62]. For both cases, we then compute the free energy difference between potentials $U(S;\hat{\theta})$ and $U(S;\theta_{i})$ using the differentiable free-energy perturbation relation [60]
@@
ReSolv outperforms classical explicit solvent force fields on FreeSolv database
In this work, we employ ReSolv to learn the hydration free energy of small organic molecules, which, by construction, also yields an ML potential for molecules in implicit water. The architecture of the ML potential is based on NequIP [63], a data efficient E(3)-equivariant graph neural network. For stage one of training, we use the QM7-X dataset [54], providing the target DFT energies and forces. The dataset consists of 4.2 million small organic molecule samples, including equilibrium states, structural isomers, structural stereoisomers, and off-equilibrium configurations. The considered molecules are composed of the heavy atom set {C, N, O, S, Cl}. They have up to seven heavy atoms or four to twenty-three atoms in total, including hydrogen. We randomly split the dataset into train (89.8%), validation (10.0%), and test (0.2%) sets. The ReSolv vacuum model ($U_{\mathrm{vac}}$) achieves a test set mean absolute error (MAE) within literature benchmarks [64, 65] (Table 1).
@@
Table 1: Top: Mean absolute error (MAE) of energy and force predictions computed on 10100 random test samples in the QM7-X dataset. Results marked with an asterisk (*) were computed by Duval et al. [64] on a different test set. Bottom: The MAE and root mean square error (RMSE) of hydration free energy prediction of different models for the same test set molecules in the FreeSolv database. The best-performing models are highlighted in bold.
@@
For stage two of training, we utilize the FreeSolv database [14]. It contains hydration free energies of 643 small molecules with neutral charges, representing compounds relevant to drug-like molecules. Keeping consistent with the heavy atom types of the QM7-X dataset, i.e., {C, N, O, S, Cl}, we extract the corresponding subset with 559 molecules. We also excluded seventeen molecules for which the ReSolv vacuum model yielded unstable simulations and an additional five molecules that were unstable during training. We split this subgroup into train and test sets of roughly 70% and 30% proportions, which amounts to 375 and 162 molecules, respectively. When dividing the data, we kept the occurrences of various heavy atom combinations consistent within the two sets, i.e., the training set contains roughly 70% of the molecules with heavy atom combinations N, N-C, N-C-O, etc. This choice ensures that the various heavy atom combinations are represented during training and testing. In addition, we ensured that the test set consists exclusively of molecules with chemical functional groups that appear in the training set, i.e., any molecule that is a lone representative of a functional group formed part of the training set. The functional group classification was taken from the FreeSolv database [13]. The polyfunctional molecules were classified according to the first listed functional group as in Ref. [31]. Apart from satisfying these two conditions, the molecules are split randomly.
@@
The ReSolv’s test set errors are reported in Table 1 and compared with the explicit solvent classical atomistic models Amber (GAFF) and CHARMM (CGenFF). We compute the errors for the same test set molecules to enable a direct comparison. Data for Amber is provided in the FreeSolv database [13], while for CHARMM, the values are taken from Ref. [31]. In the CHARMM error computation, two molecules (Mobley ID 6359135 and 2146331) are excluded because data is not provided in Ref. [31] since the force field parameters could not be generated. The ReSolv’s MAE is close to the average uncertainty of the FreeSolv database (0.6 kcal/mol). Moreover, ReSolv outperforms the explicit solvent classical atomistic models by a large margin. Amber and CHARMM models display similar MAE/RMSE for the molecules in the training set (Supplementary Figure 1), confirming that the test set is representative of the entire considered database. The comparison with the Amber and CHARMM models is somewhat unfair, given that the two models were not explicitly parametrized to reproduce the hydration free energy data. Nevertheless, reparametrizations and corrections in this direction were previously attempted, and the resulting models still scored lower than ReSolv. For example, Boulanger et al. rescaled the Amber (GAFF) molecule-water van der Waals dispersion interaction to better reproduce the hydration free energy and reported an MAE of 0.79 kcal/mol for the final optimized model [30]. In another example, Scheen et al. trained an ML model to correct the Amber (GAFF) force field predictions [28]. The hybrid Amber/ML approach achieved an MAE of 0.76 kcal/mol on the SAMPL4 test data containing 47 samples of the FreeSolv database.
@@
To further demonstrate this point we plot the error distributions relative to the experimental values (Fig. 2b). The distributions are skewed to the left for classical force fields, indicating that the majority of predictions are overestimated. The Amber model, for example, overestimates the hydration free energies on average by more than 1 kcal/mol as already reported by Boulanger et al. [30]. Conversely, no systematic errors are found for the ReSolv model, evidenced by the symmetric distribution about zero. These findings are reflected in the percentage of predictions within the experimental uncertainty (Fig. 2c). The ReSolv model scores the highest, followed by CHARMM and Amber in last place, albeit Amber exhibiting lower MAE and RMSE than the CHARMM model.

### v2

@@ Abstract
Machine learning (ML) potentials are a powerful tool in molecular modeling, enabling ab initio accuracy for comparably small computational costs. Nevertheless, all-atom simulations employing best-performing graph neural network architectures are still too expensive for applications requiring extensive sampling, such as free energy computations. Implicit solvent models could provide the necessary speed-up due to reduced degrees of freedom and faster dynamics. Here, we introduce a Solvation Free Energy Path Reweighting (ReSolv) framework to parametrize an implicit solvent ML potential for small organic molecules that accurately predicts the hydration free energy, an essential parameter in drug design and pollutant modeling. With a combination of top-down (experimental hydration free energy data) and bottom-up (ab initio data of molecules in a vacuum) learning, ReSolv bypasses the need for intractable ab initio data of molecules in explicit bulk solvent and does not have to resort to less accurate data-generating models. On the FreeSolv dataset, ReSolv achieves a mean absolute error close to average experimental uncertainty, significantly outperforming standard explicit solvent force fields. Compared to the explicit solvent ML potential, ReSolv offers a computational speedup of four orders of magnitude and attains closer agreement with experiments. The presented framework paves the way toward deep molecular models that are more accurate yet computationally cheaper than classical atomistic models.
preprint: AIP/123-QED
I Introduction
@@
Solvation free energy, and notably hydration free energy, is generally recognized as a fundamental thermodynamic quantity of interest in computational chemistry. Defined as the work done when transferring a molecule from the gas phase to the solution (water in the case of hydration free energy), it enables the computation of several key physicochemical properties of molecules, such as solubility, partition coefficients, activity coefficients, and binding free energies in solutions Fossat, Zeng, and Pappu (2021); Nerenberg et al. (2012). These properties are of great importance to the pharmaceutical, environmental, and materials sciences Ratkova, Palmer, and Fedorov (2015); Boobier et al. (2020); Deng and Jia (2020); Zubatyuk et al. (2019); Rauer and Bereau (2020); Hutchinson and Kobayashi (2019); Lim and Jung (2019), prompting the organization of blind prediction SAMPL challenges Nicholls et al. (2008); Geballe et al. (2010); Guthrie (2014) with hydration free energy as one of the main targets. In addition, Mobley et al. compiled and curated a FreeSolv database of experimentally measured hydration free energies for small neutral molecules in water Mobley and Guthrie (2014); Duarte Ramos Matos et al. (2017).
@@
A wide spectrum of methods is available to calculate solvation free energy, ranging from traditional approaches such as continuum solvation models Klamt and Diedenhofen (2010); Ehlert et al. (2021) to recent machine learning (ML) algorithms Riniker (2017); Zhang et al. (2023); Lim and Jung (2021); Low, Coote, and Izgorodina (2022); Zhang, Xia, and Zhang (2022); Wu et al. (2018); Yang et al. (2019); Cho and Choi (2019); Pathak, Mehta, and Priyakumar (2021); Chen et al. (2021a) and their combinations Alibakhshi and Hartke (2021); Scheen et al. (2020); Weinreich, Browning, and von Lilienfeld (2021). The alchemical methods with Molecular Dynamics (MD) simulations Duarte Ramos Matos et al. (2017); Boulanger et al. (2018); Karwounopoulos et al. (2023) are typically assumed to be highly accurate but computationally expensive Luukkonen et al. (2020); Zhang et al. (2017). However, both the fidelity and the efficiency highly depend on the explicitly treated degrees of freedom and the employed potential energy model.
@@
In implicit solvent models, the solvent molecules (e.g., water molecules) are not explicitly present in the system (as in explicit solvent models); instead, the interactions are modified to account for the solvent effects Onufriev (2008). The number of degrees of freedom is thereby greatly reduced, resulting in large computational gains. Classical implicit models treat the solvent as a continuum medium with specific dielectric and interfacial properties. The typical approach decouples the electrostatic (polar) and nonpolar interactions. The former can be approximated by solving the Poisson–Boltzmann equation or further simplified with the popular generalized Born model, while the latter is most often estimated via the solvent-accessible surface area Onufriev (2008); Roux and Simonson (1999). While recent advances, such as using ML to predict the generalized Born radii Mahmoud et al. (2020), have increased the accuracy of these models, they are still in considerable disagreement with experimental data and explicit solvent models Zhou and Berne (2002); Chen et al. (2021a); Cumberworth, Bui, and Gsponer (2016). In particular, the solvation free energy root mean square error (RMSE) is approximately 3.6 kcal/mol for both the Poisson-Boltzmann and Generalized Born models. By specifically optimizing the nonpolar interactions using hydration free energy data, the test set RMSE can be decreased to 1.68 kcal/mol Brieg et al. (2017).
Higher accuracy can be achieved through ab initio methods, based on quantum chemical calculations, albeit at the cost of computational demand. Commonly utilized is the COSMO family Klamt and Schüürmann (1993); Klamt and Eckert (2000), with COSMO-RS (Conductor-like Screening Model for Real Solvents) being particularly notable in property predictions. In particular, extensions thereof have achieved high fidelity predictions of solvation free energies (MAE=0.52 kcal/mol)Klamt and Diedenhofen (2015).
@@
An alternative to the computational hurdle imposed by ab initio methods, without significantly compromising predictive accuracy, is offered by the many-body, flexible potential energy surfaces, that characterize ML potentials. During the past decade, ML potentials were deployed to derive atomistic Friederich et al. (2021) and coarse-grained Gkeka et al. (2020) models based on training data provided by high-fidelity simulations Thaler, Stupp, and Zavadlav (2022), experiments Thaler and Zavadlav (2021), or both Röcken and Zavadlav (2024). Deriving empirical force fields from experimental data has been extensively studied, see e.g. Refs. Ding (2024); Polêto and Lemkul (2022); Cesari et al. (2019); Fröhlking et al. (2020), and this approach naturally extends to ML potentials. In the context of implicit water models, the capacity of deep ML potentials is large enough to compensate for the removal of solvent degrees of freedom. The predicted solute properties, such as the conformational landscape of proteins, match those obtained from the reference explicit water model Vlachas et al. (2021); Coste et al. (2023); Chen et al. (2021a); Wang et al. (2021); Wang, Axelrod, and Gómez-Bombarelli (2020); Husic et al. (2020); Wang and Gómez-Bombarelli (2019); Yao et al. (2023). However, these studies used classical atomistic force fields as the data-generating model, inherently limiting the attainable accuracy of the resulting ML potential. As shown previously, and also in this work, classical atomistic models such as General Amber Force Field (GAFF) and CHARMM General Force Field (CGenFF) systematically overestimate hydration free energies Boulanger et al. (2018). Modifying the Lennard-Jones parameters yields an improved agreement with experiments but can, at the same time, negatively affect other properties Boulanger et al. (2018).
@@
On the other hand, more accurate ab initio training data is prohibitively expensive for solutes in explicit bulk water. Consequently, frequently used Density Functional Theory (DFT) training databases such as QM9 Ramakrishnan et al. (2014), ANI Smith, Isayev, and Roitberg (2017), and QM7-X Hoja et al. (2021) contain small organic molecules in vacuum or at best contain samples with a few randomly placed water molecules around solutes Unke and Meuwly (2019); Basdogan et al. (2019). Moreover, implicit solvent models are coarse-grained models. When trained with the common force-matching approach Noid et al. (2008), coarse-grained ML models require much more data than their atomistic counterparts Durumeric et al. (2023). The underlying reason is the surjective atomistic-to-coarse-grained mapping, resulting in noisy force labels. As an illustration, parametrization of coarse-grained models for alanine dipeptide in implicit water required a dataset with $10^{6}$ configurations Chen et al. (2021b); Thaler, Stupp, and Zavadlav (2022), a number easily obtainable with classical force fields but vastly out of reach for ab initio calculations.
@@
II Methods
II.1 Solvation Free Energy Path Reweighting (ReSolv)
@@
The top-down training on experimental data is not as straightforward as bottom-up training in the first stage because $\Delta A$ is not an output of the ML model but rather evaluated from an MD simulation driven by an ML potential. We employ a variation of the Differential Trajectory Reweighting (DiffTRe) method Thaler and Zavadlav (2021) that avoids exploding gradients and reduces the computational and memory requirements compared to alternative gradient computation via backpropagation through the MD simulation. Before continuing, we first give a brief summary of the DiffTRe method.
Consider the task of matching a time-independent observable $O$, e.g., by postulating the following loss function $L=(\langle O\rangle_{\theta}-O_{\mathrm{exp}})^{2}$, where $\langle\rangle_{\theta}$ denotes the ensemble average with respect to the canonical distribution using ML potential $U_{\theta}=U(S;\theta)$. DiffTRe leverages thermodynamic perturbation theory Zwanzig (1954), stating that $\langle O\rangle_{\theta}$ can be estimated from $N$ decorrelated states $\{S_{j}\}_{\hat{\theta}}$ generated by a reference potential $U_{\hat{\theta}}=U(S;\hat{\theta})$, that is
@@
where $\beta=1/(k_{B}T)$, $k_{B}$ is Boltzmann’s constant and $T$ temperature. Due to limited sampling, the estimation should only be used if the states generated by the reference potential $U_{\hat{\theta}}$ are statistically close to states that would have been sampled from the potential $U_{\theta}$. The distribution overlap is captured with the effective sample size Carmichael and Shell (2012) given by
@@
Here, we instead use the DiffTRe learning process itself as a free energy integration path. We initialize the parameters with the pre-trained $\theta_{\mathrm{vac}}$ parameters and iteratively perturb them during training towards the desired $\theta_{\mathrm{sol}}$ parameters (see Fig. 1). At each update step, we compute the free energy difference between the new and previous potential using a hybrid of the free energy perturbation Zwanzig (1954) and the Bennett Acceptance Ratio (BAR) methods Bennett (1976). The free energy differences are accumulated along the training, which, at the end of the training, yield the total free energy difference between $U_{\mathrm{vac}}$ and $U_{\mathrm{sol}}$ or $\Delta A$.
Explicitly, let us consider the update step from $\theta_{i}$ to $\theta_{i+1}$. At this point, due to reweighting, the reference trajectory may have been generated at an earlier step $i-k$ with the ML potential $U(S;\theta_{i-k})$, where $k\leq i$. Since this is the reference trajectory, $\hat{\theta}=\theta_{i-k}$. As above, we denote the states generated by the reference potential with $\{S_{j}\}_{\hat{\theta}}$. The corresponding free energy difference is $\Delta A=\Delta A_{\theta_{0}\to\theta_{i-k}}$, which was already computed at step $i-k$. Firstly, we compute $N_{\mathrm{eff}}$ (Eq. 3). If the reweighting criterion is satisfied, the trajectory is reused. In the opposite case, we generate a new reference trajectory with the current ML potential $U(S;\theta_{i})$. Thus, $\hat{\theta}=\theta_{i}$. We also update $\Delta A=\Delta A+\Delta A^{BAR}_{\theta_{i-k}\to\theta_{i}}$, where $\Delta A^{BAR}_{\theta_{i-k}\to\theta_{i}}$ denotes the free energy difference between steps $i-k$ and $i$ estimated with the BAR method Bennett (1976). For both cases, we then compute the free energy difference between potentials $U(S;\hat{\theta})$ and $U(S;\theta_{i})$ using the differentiable free-energy perturbation relation Zwanzig (1954)
@@
II.2 Model training and validation
Both vacuum and solvent ML potentials are based on the GNN architecture NequIP Batzner et al. (2022) as implemented in JAX MD Schoenholz and Cubuk (2020). Computations were performed with double precision on Nvidia A100 80GB GPUs. The architectural hyperparameters of NequIP are listed in supplementary material Table S2.
where $\gamma_{U}$ and $\gamma_{F}$ are weighting factors, $N_{bs}$ is the batch size, $N_{i,a}$ the number of atoms of molecule $i$, and $k$ iterates over the x-, y-, z-dimensions. $U_{i}$ is the energy of the $i$-th molecule in the batch, and $F_{ijk}$ is the force in direction k of atom j from the i-th molecule in the batch. Subscript DFT denotes the ab initio data target. We normalize the target energies and forces. The energies are shifted by the mean energy of the training data and scaled with the average root mean squared force of the training data. The forces are scaled with the average root mean square force of the training data. Further, we do not employ learnable scaling or shifting and set the per-atom scaling to 1 and the per-atom shift to 0. To pick the best model, we employ early stopping on the validation dataset. For the numerical optimization hyperparameters, see supplementary material Table S3.
In the second stage, we employ the RDKit rdk to generate the 3D structure of a molecule based on the SMILES string provided by the FreeSolv database. Next, we perform the energy minimization of structures with the MMFM94 force field Halgren (1996a, b, c); Halgren and Nachbar (1996); Halgren (1996d). The obtained configurations are used to run a 300 ps initial equilibration simulation followed by a 200 ps production run for each molecule. All simulations are performed in the NVT ensemble and numerically integrated with the velocity Verlet scheme using a 1 fs timestep. To match the conditions of the experimental data the temperature is kept at 298.15 K with the Langevin thermostat with damping factor set to 1/ps. We check the stability of our simulations by evaluating that the last configuration of the molecules can be represented by a single graph and that no atoms diverged from the molecule. Additionally, to see whether our models yield physically reasonable trajectories, we compare the bonds, angles, and dihedrals of the $U_{sol}$ trajectories for three randomly selected molecules to the bonds, angles, and dihedrals as obtained after an energy minimization with the UFF force field using RDKit rdk ; see supplementary material Fig. S9. In training, new simulations are initialized with the last configuration of the previous simulation and consist of 50 ps equilibration and a 200 ps production run. We sample every 5 ps during the production run, i.e., 40 samples per trajectory. The reweighting effective sample size threshold is fixed to $\bar{N}_{\mathrm{eff}}=0.9$. The numerical optimization hyperparameters are reported in supplementary material Table S3.
The reported hydration free energies are computed with the BAR method using only the end states, i.e., vacuum and water states. As previously reported, intermediate states are not necessary for implicit solvent models due to the sufficient overlap between the vacuum and solvated ensembles of the solute Zhang et al. (2017). We first run 300 ps of equilibration and sample every 5 ps in the subsequent 200 ps production run. We perform convergence tests (supplementary material Fig. S8) for 12 randomly selected molecules in the test set to ensure sufficient sampling.
III Results and discussion
III.1 ReSolv outperforms classical explicit solvent force fields on FreeSolv database
In this work, we employ ReSolv to learn the hydration free energy of small organic molecules, which, by construction, also yields an ML potential for molecules in implicit water. The architecture of the ML potential is based on NequIP Batzner et al. (2022), a data efficient E(3)-equivariant graph neural network. For stage one of training, we use the QM7-X dataset Hoja et al. (2021), providing the target DFT energies and forces. The dataset consists of 4.2 million small organic molecule samples, including equilibrium states, structural isomers, structural stereoisomers, and off-equilibrium configurations. The considered molecules are composed of the heavy atom set {C, N, O, S, Cl}. They have up to seven heavy atoms or four to twenty-three atoms in total, including hydrogen. We randomly split the dataset into train (89.8%), validation (10.0%), and test (0.2%) sets. The ReSolv vacuum model ($U_{\mathrm{vac}}$) achieves a test set mean absolute error (MAE) within literature benchmarks Duval et al. (2023); Unke et al. (2021) (Table 1).
@@ III.1 ReSolv outperforms classical explicit solvent force fields on FreeSolv database
Table 1: Top: Mean absolute error (MAE) of energy and force predictions computed on 10100 random test samples in the QM7-X dataset. Results marked with an asterisk (*) were computed by Duval et al. Duval et al. (2023) on a different test set. Bottom: The MAE and root mean square error (RMSE) of hydration free energy prediction of different models for the same test set molecules in the FreeSolv database. The best-performing models are highlighted in bold.
@@
For stage two of training, we utilize the FreeSolv database Duarte Ramos Matos et al. (2017). It contains hydration free energies of 643 small molecules with neutral charges, representing compounds relevant to drug-like molecules. Keeping consistent with the heavy atom types of the QM7-X dataset, i.e., {C, N, O, S, Cl}, we extract the corresponding subset with 559 molecules. We also excluded seventeen molecules for which the ReSolv vacuum model yielded unstable simulations and an additional five molecules that were unstable during training. We split this subgroup into train and test sets of roughly 70% and 30% proportions, which amounts to 375 and 162 molecules, respectively. When dividing the data, we kept the occurrences of various heavy atom combinations consistent within the two sets, i.e., the training set contains roughly 70% of the molecules with heavy atom combinations N, N-C, N-C-O, etc. This choice ensures that the various heavy atom combinations are represented during training and testing. In addition, we ensured that the test set consists exclusively of molecules with chemical functional groups that appear in the training set, i.e., any molecule that is a lone representative of a functional group formed part of the training set. The functional group classification was taken from the FreeSolv database Mobley and Guthrie (2014). The polyfunctional molecules were classified according to the first listed functional group as in Ref. Karwounopoulos et al. (2023). Apart from satisfying these two conditions, the molecules are split randomly.
@@
The ReSolv’s test set errors are reported in Table 1 and compared with the explicit solvent classical atomistic models Amber (GAFF) and CHARMM (CGenFF). We compute the errors for the same test set molecules to enable a direct comparison. Data for Amber is provided in the FreeSolv database Mobley and Guthrie (2014), while for CHARMM, the values are taken from Ref. Karwounopoulos et al. (2023). In the CHARMM error computation, two molecules (Mobley ID 6359135 and 2146331) are excluded because data is not provided in Ref. Karwounopoulos et al. (2023) since the force field parameters could not be generated. The ReSolv’s MAE is close to the average uncertainty of the FreeSolv database (0.6 kcal/mol). Moreover, ReSolv outperforms the explicit solvent classical atomistic models by a large margin. Amber and CHARMM models display similar MAE/RMSE for the molecules in the training set (supplementary material Fig. S1), confirming that the test set is representative of the entire considered database. In addition, we trained the ReSolv model using a random splitting of the data and achieved similar performance. In particular, the obtained MAE and RMSE are 0.65 and 0.91 kcal/mol, respectively. The corresponding errors for the classical force fields remain substantially higher (supplementary material Fig. S2). The comparison with the Amber and CHARMM models is somewhat unfair, given that the two models were not explicitly parametrized to reproduce the hydration free energy data. Nevertheless, reparametrizations and corrections in this direction were previously attempted, and the resulting models still scored lower than ReSolv. For example, Boulanger et al. rescaled the Amber (GAFF) molecule-water van der Waals dispersion interaction to better reproduce the hydration free energy and reported an MAE of 0.79 kcal/mol for the final optimized model Boulanger et al. (2018). In another example, Scheen et al. trained an ML model to correct the Amber (GAFF) force field predictions Scheen et al. (2020). The hybrid Amber/ML approach achieved an MAE of 0.76 kcal/mol on the SAMPL4 test data containing 47 samples of the FreeSolv database.
@@
To further demonstrate this point we plot the error distributions relative to the experimental values (Fig. 2b). The distributions are skewed to the left for classical force fields, indicating that the majority of predictions are overestimated. The Amber model, for example, overestimates the hydration free energies on average by more than 1 kcal/mol as already reported by Boulanger et al. Boulanger et al. (2018). Conversely, no systematic errors are found for the ReSolv model, evidenced by the symmetric distribution about zero. These findings are reflected in the percentage of predictions within the experimental uncertainty (Fig. 2c). The ReSolv model scores the highest, followed by CHARMM and Amber in last place, albeit Amber exhibiting lower MAE and RMSE than the CHARMM model.